OpenStudy (lilkg77):

A principal of $400 is invested in an account at 6% per year compounded annually. What is the total amount of money in the account after 5 years? A. $535.29 B. $526.00 C. $520.00 D. $531.37

OpenStudy (anonymous):

This can be expressed as the function \[y=400(1.06^x)\] Or you could do it by taking 6% of 400, adding it to 400, taking 6% of the new number, adding that to the new number, etc.

jhonyy9 (jhonyy9):

first can you calculi 6% of 400 = ? - so and this just multiplie by 5 assum to 400 and will get the right total amount after 5 year

OpenStudy (anonymous):

Where y is the money after a period of time and x is the number of years

jhonyy9 (jhonyy9):

@lilkg77 do you understand it ?

OpenStudy (lilkg77):

Yes a little @jhonyy9

OpenStudy (anonymous):

Well the way I got that expression is exponential equations are always written as c(t^x) Where c is the initial and t is the rate of change I put 1.06 instead of .06 for t because it is growth not decay in this case

OpenStudy (dayakar):

p=$400 , r = 6% , n= 5 \[Amount (A) = p ( 1+ \frac{ r }{ 100 })^{n}\] substitute the values simplify it

OpenStudy (wolf1728):

If it is annual compounding the formula is just Total = Principal × ( 1 + Rate ) ^ years

OpenStudy (wolf1728):

Total = 400 × ( 1.06 )^5

OpenStudy (lilkg77):

Thanks @wolf1728

OpenStudy (wolf1728):

u r welcome lilkg77 :-)